A national bank published a report on improvements in customer satisfaction and loyalty. the report concluded that 48% of the bank customers expressed customer delight regarding the banks customer services. suppose you surveyed 350 custimers in your town to inquire about customer delight towards this bank.
a)find the probability that the majority of the people in your sample of 350 customers would express customer delight.
b)the probability of 70% that the sample proportion of respondents expressing customer delight will be less than what percentage?
c)the probability is 80% that the sample proportion will be contained within what symmtetrical limits of the population proportion

Answer by stanbon(75887)

A national bank published a report on improvements in customer satisfaction and loyalty. the report concluded that 48% of the bank customers expressed customer delight regarding the banks customer services. suppose you surveyed 350 customers in your town to inquire about customer delight towards this bank.
----
mean of the sample means = 0.48
std of the sample means = sqrt(0.48*0.52/350) = 0.0267
----------------------------------------
a)find the probability that the majority of the people in your sample of 350 customers would express customer delight.:::
z(0.5) = (0.5-0.48)/0.0267) = 0.7491
P(p-hat > 0.5) = P(z > 0.7491) = 0.2269
----------------------------------
b)the probability of 70% that the sample proportion of respondents expressing customer delight will be less than what percentage?
invNorm(0.7) = 0.5244
p-hat = zs+u
p-hat = 0.5244*0.0267+0.48 = 0.4940
--------------------------------------------
c)the probability is 80% that the sample proportion will be contained within what symmtetrical limits of the population proportion
invNom(0.1) = -1.2816
lower limit = p-hat = -1.2816*0.0267+0.48 = 0.4458
upper limit = p-hat = +1.2816*0.0267+0.48 = 0.5142
======================================================
Cheers,
Stan H.
=============